A current loop contains 100 turns of wire and has dimensions of 0.020m by 0.030m. The current through the loop is 0.10A and the external magnetic field is directed to produce maximum torque. If the magnitude of the magnetic field is 0.030T, determine the a) torque on the loop and b) magnetic moment of coil.
Answer #1

In case of coil torque induced in the coil is given as \tau= NIBA

Here N = 100

A= 6× 10-4m^2

B=0.030 T

I= 0.10

So \tau = 18×10-5 Joule

Magnetic moment is given as

M= NIA

M= 6× 10-4 Ampere m​​​​​​2